1, A 0.30 m radius automobile tire accelerates from rest at a constant angular acceleration of 0.758 rad/s2 for 2.23 s. What is its final angular speed in rad/s?

2. A turntable has a moment of inertia of 3.00 x 10-2 kg-m2 and spins freely on a frictionless bearing at 46 rev/min. A 1.11 kg ball of putty is dropped vertically onto the turntable and sticks at a point 0.100m from the center. What is the new rate of rotation of the system in rev/min?
Answer #1

Question 1

Radius of the automobile tire = R = 0.3 m

Initial angular speed of the tire = 1 = 0 rad/s (At rest)

Final angular speed of the tire = 2

Angular acceleration of the tire = = 0.758 rad/s2

Time period of acceleration = T = 2.23 sec

2 = 1 + T

2 = 0 + (0.758)(2.23)

2 = 1.69 rad/s

Final angular speed of the tire = 1.69 rad/s

Question 2

Moment of inertia of the turntable = I = 3 x 10-2 kg.m2

Initial angular speed of the turn table = 1 = 46 rev/min

Converting from rev/min to rad/s

1 = 46 x (2/60) rad/s

1 = 4.817 rad/s

Mass of the ball of putty = M = 1.11 kg

Distance of the point where the putty sticks from the center of the turntable = R = 0.1 m

New angular speed of the system = 2

By conservation of angular momentum,

I1 = (I + MR2)2

(3x10-2)(4.817) = (3x10-2 + (1.11)(0.1)2)2

2 = 3.516 rad/s

Converting to rev/min,

2 = 3.516 x (60/2) rev/min

2 = 33.58 rev/min

New rate of rotation of the system = 33.58 rev/min